Which exerts a stronger gravitational pull on Earth's Oceans, sun or moon? Explain/Prove it. (You may assume physical constants like Rearth, Mearth, G, etc.)
Answer #1

here,

the distance of Earth from moon , r1 = 3.844 * 10^8 m

mass of moon , M1 = 7.34 * 10^22 kg

mass of Earth , Me = 5.98 * 10^24 kg

the gravitational pull on Earth due to Moon , F1 = G * M1 * Me /r1^2

F1 = 6.67 * 10^-11 * 7.34 * 10^22 * 5.98 * 10^24 /(3.844 * 10^8)^2

F1 = 1.98 * 10^20 N

the distance of Earth from Sun , r2 = 1.47 * 10^11 m

mass of Sun , M2 = 2 * 10^30 kg

mass of Earth , Me = 5.98 * 10^24 kg

the gravitational pull on Earth due to Sun , F2 = G * M2 * Me /r2^2

F2 = 6.67 * 10^-11 * 2 * 10^30 * 5.98 * 10^24 /(1.47 * 10^11)^2

F2 = 3.69 * 10^22 N

so, the Sun exerts stronger pull than Moon on Earth